### v1

Abstract
This paper is concerned with the approximation of probability distributions known up to normalization constants, with a focus on Bayesian inference for large-scale inverse problems in scientific computing. In this context, key challenges include costly repeated evaluations of forward models, multimodality, and inaccessible gradients for the forward model. To address them, we develop a variational inference framework that combines Fisher-Rao natural gradient with specialized quadrature rules to enable derivative free updates of Gaussian mixture variational families. The resulting method, termed Derivative Free Gaussian Mixture Variational Inference (DF-GMVI), guarantees covariance positivity and affine invariance, offering a stable and efficient framework for approximating complex posterior distributions. The effectiveness of DF-GMVI is demonstrated through numerical experiments on challenging scenarios, including distributions with multiple modes, infinitely many modes, and curved modes in spaces with up to hundreds of dimensions. The method’s practicality is further demonstrated in a large-scale application, where it successfully recovers the initial conditions of the Navier-Stokes equations from solution data at positive times.
keywords:
Bayesian Inverse Problems, Variational Inference, Derivative Free Methods, Multimodal, Gaussian Mixtures.
{AMS}
68Q25, 68R10, 68U05
1 Introduction
Sampling a target probability distribution known up to normalization constants is a classical problem in scientific computing. Specifically, in Bayesian inverse problems [39, 57], the goal is to recover an unknown parameter $\theta\in\mathbb{R}^{N_{\theta}}$ from noisy observation $y\in\mathbb{R}^{N_{y}}$, through the equation
(1) $y=\mathcal{G}(\theta)+\eta.$
Here, $\mathcal{G}$ represents a forward map which, for the problems we consider, is nonlinear and requires solving a partial differential equation (PDE) for each evaluation. The observational noise $\eta$ follows a Gaussian distribution: $\eta\sim\mathcal{N}(0,\Sigma_{\eta})$. Within the Bayesian framework, we assign a Gaussian prior $\mathcal{N}(r_{0},\Sigma_{0})$ to the unknown parameter $\theta$, resulting in a posterior distribution from which we aim to draw samples
(2) $\rho_{\rm post}(\theta)$ $\propto\exp(-\Phi_{R}(\theta)),\ \Phi_{R}(\theta)=\frac{1}{2}\lVert\Sigma_{\eta}^{-\frac{1}{2}}(y-\mathcal{G}(\theta))\rVert^{2}+\frac{1}{2}\lVert\Sigma_{0}^{-\frac{1}{2}}(\theta-r_{0})\rVert^{2}.$
It is worth noting that in this context, $\Phi_{R}$ adopt a nonlinear least-squares structure, with an augmented map $\mathcal{F}(\theta)$ satisfying
(3) $\Phi_{R}(\theta)=\frac{1}{2}\mathcal{F}(\theta)^{T}\mathcal{F}(\theta),\ \ \mathcal{F}(\theta)=\begin{bmatrix}\Sigma_{\eta}^{-\frac{1}{2}}\bigl{(}y-\mathcal{G}(\theta)\bigr{)}\\ \Sigma_{0}^{-\frac{1}{2}}\bigl{(}r_{0}-\theta\bigr{)}\end{bmatrix}.$
The primary goal of this paper is to develop an efficient approach for the approximate sampling of such $\rho_{\rm post}(\theta)$.
1.1 Challenges
For many Bayesian inverse problems in scientific applications, computing gradients of $\Phi_{R}$ might be infeasible, as it requires derivatives of large-scale PDE-based models $\mathcal{G}$ that may be black-box (e.g., climate models [55, 54]), use non-differentiable numerical methods (e.g., embedded boundary methods [51, 31, 30, 9] and adaptive mesh refinement [3, 6]), or model discontinuous physics (e.g., in fracture [48] or cloud modeling [58, 44]). Derivative free methods are thus needed; see a review of methodologies in Section 1.3.2. While there are a few Markov Chain Monte Carlo (MCMC) and Sequential Monte Carlo (SMC) approaches that do not require gradients, they often require numerous function evaluations, particularly in high dimensions, to achieve convergence or mitigate weight collapse. This makes them impractical given the high computational cost of evaluating $\mathcal{G}$. Furthermore, the multimodality of $\rho_{\rm post}(\theta)$ causes MCMC methods to struggle with mode transitions [23, 24]. We note that missing modes can lead to significant prediction errors in scientific applications [59].
Variational inference offers a promising alternative, with black-box variational inference (BBVI) [52] being a popular approach for achieving derivative free sampling. BBVI typically relies on stochastic approximations to estimate gradients which exist high variance. This often makes BBVI unreliable, requiring variance reduction techniques and extensive time step tuning. Even with these improvements, small time steps are necessary to maintain stability (see Section 5.2). Kalman methodology is also used to develop derivative-free variational inference methods [14, 32]. While these approaches can be effective for posterior distributions close to Gaussian or Gaussian mixture with separated components, they often become unstable in the presence of non-Gaussian or strongly curved modes (see [14, Appendix E]).
In this work, we explore guidelines for designing stable derivative free Gaussian mixture variational inference methods. Building on these guidelines, we propose a novel approach that requires no hyperparameter tuning while achieving both high accuracy and computational efficiency.
1.2 Contributions
Specifically, our contributions are as follows:
  1. 1.
    We propose derivative free quadrature rules for approximating the expectations of $\Phi_{R}$ and Gaussian mixtures, their gradients, and their Hessians with respect to any Gaussian density. These methods are exact for linear $\mathcal{F}$ implying low variance, and capture as much curvature information of $\mathcal{F}$ as possible. Moreover, the computational cost, in terms of evaluating the forward mapping $\mathcal{F}(\theta)$, scales linearly with the dimensionality of $\theta$.
  2. 2.
    Building on these quadrature rules, we develop a derivative free variational inference approach that combines Fisher-Rao natural gradient and Gaussian mixture approximations (DF-GMVI). The DF-GMVI method ensures covariance positivity and exhibits affine invariant properties, which together contribute to its superior stability even when we use large time steps. These guidelines can also be applied to improve the robustness of other Gaussian mixture variational inference methods.
  3. 3.
    We demonstrate that DF-GMVI effectively captures multiple or even infinite modes, as well as curved modes, in model problems involving up to hundreds of dimensions. Additionally, it performs well in PDE-based applications, such as reconstructing the Navier-Stokes initial condition from solution data at positive times.
1.3 Literature Review
The review of relevant literature concerns variational inference, and related derivative free sampling approaches.
1.3.1 Variational Inference
The main idea of variational inference is to approximate a target density $\rho_{\rm post}$ within a variational family of densities $Q$ from the view of optimization, that is, to identify the member of this family that minimizes an energy function $\mathcal{E}$:
$\rho=\operatorname*{arg\,min}_{\rho\in Q}\mathcal{E}(\rho,\rho_{\rm post}),$
where the minimizer coincides with $\rho_{\rm post}$. The energy function is often chosen as the Kullback-Leibler divergence (4). Along the gradient flow of the energy function, $\frac{\mathrm{d}\rho}{\mathrm{d}t}=-\nabla_{M}\mathcal{E}(\rho,\rho_{\rm post})$, with respect to the metric $M$, the density $\rho$ gradually converges towards the target density. In practice, variational inference methods can be categorized into two categories: non-parametric and parametric approaches.
In non-parametric variational inference, the variational distribution is represented by an interacting particle system. The choice of metrics [13] determines the gradient flow, including the Wasserstein gradient flow [35, 12, 40], Fisher-Rao gradient flow [47, 17, 11, 63], Wasserstein-Fisher-Rao gradient flow [45], Kalman-Wasserstein gradient flow [22] and Stein gradient flow [42]. The interacting particle system evolves according to the gradient flow, gradually approximating the target distribution. The convergence rate and accuracy depend on the number of particles used to represent the distribution and the quality of their representation.
In parametric variational inference, the variational densities used to approximate the target density are parametrized. A common choice for the variational family is the Gaussian, which leads to Gaussian variational inference [50, 40, 5]. To account for the geometric structure of parameters, the gradient is preconditioned by the Fisher information matrix, resulting in the Fisher-Rao natural gradient [1, 50]. This approach leverages Fisher information matrix as a substitute for the Hessian, thereby outperforming standard gradient descent [46]. Gaussian approximations have also been explored for Stein gradient flows [43] and Wasserstein gradient flows [40, 16]. However, the Gaussian approximation has limited expressive power, prompting the use of more flexible variational families, such as Gaussian mixture. These include Gaussian mixture approximation of the natural gradient flow [41, 14] and Wasserstein gradient flow [40, 34]. The present work focuses on Gaussian mixture approximation of the natural gradient flow, with an emphasis on stable and efficient derivative free approximations.
1.3.2 Derivative Free Sampling Approaches
A large class of derivative free sampling approaches are Markov chain Monte Carlo (MCMC) methods [26, 25], where a derivative free proposal is used to move particles. The main challenges of these methods are the absence of stopping criteria and slow convergence, which worsen as the dimensionality of the problem increases. Several improvements have been proposed to address these issues, such as enhancing the proposal distribution by preserving affine invariance (e.g., the stretch move method [27]) and leveraging parallelization (e.g., with multiple chains [21, 7, 60]). This work also incorporates these two strategies.
Another class of methods is Sequential Monte Carlo (SMC)[18, 56, 4], where particles and their associated weights are updated together using importance sampling. However, resampling is necessary to address issues such as particle degeneracy, sample impoverishment, and instability[19], which arise due to the lack of gradient flow structures.
Finally, in variational inference, a major approach is black-box variational inference [52], which employs stochastic approximation via Monte Carlo methods to estimate gradients. However, the high variance of these gradient estimates often results in unstable updates and slow convergence. To address these challenges, strategies such as incorporating variance reduction techniques and adaptive learning rates to stabilize the gradient estimates have been proposed in [52, 62]. Alternatively, gradient-based variation inference methods, such as Kalman-Wasserstein gradient flow [22] and Fisher-Rao gradient flow [14], avoid direct gradient computation by leveraging Stein’s lemma: $\mathbb{E}_{\mathcal{N}(\theta;m,C)}[\nabla_{\theta}\mathcal{F}(\theta)]=C^{-1}{\rm Cov}_{\mathcal{N}(\theta;m,C)}[\mathcal{F}(\theta),\theta]$. This approach enables gradient estimation through quadrature rules originated from Kalman filtering techniques, such as the ensemble Kalman filter [20], the unscented Kalman filter [37, 36, 33], and the cubature Kalman filter [2]. Notably, the latter two achieve exactness for linear $\mathcal{F}$, indicating low variance. Despite these advancements, these approaches still require small time steps to mitigate instability, especially when the posterior is high-dimensional, exhibits multimodality, or involves complex dependencies. This work focuses on advancing quadrature rules for estimating gradients and even Hessians, which are crucial for developing stable, derivative-free variational inference methods.
1.4 Organization
The remainder of the paper is organized as follows. In Section 2, we provide an overview of natural gradient variational inference, including both Gaussian and Gaussian mixture variational families. Section 3 introduces our Derivative Free Gaussian Mixture Variational Inference (DF-GMVI), and the related theoretical insights are presented in Section 4. Numerical experiments are described in Section 5, which serve to empirically validate the theory and demonstrate the effectiveness of the proposed framework for Bayesian inference. Finally, concluding remarks are provided in Section 6.
2 Natural Gradient Variational Inference
In this section, we first briefly review natural gradient variational inference from the perspective of gradient flow. Then we discuss both the Gaussian variational family and the Gaussian mixture variational family in Sections 2.1 and 2.2, respectively. Variational inference aims to approximate the posterior distribution Eq. 2 by minimizing the Kullback–Leibler (KL) divergence [61, 5]
(4) ${\rm KL}[\rho_{a}\|\rho_{\rm post}]=\int\rho_{a}\log\Bigl{(}\frac{\rho_{a}}{\rho_{\rm post}}\Bigr{)}\mathrm{d}\theta$
over a variational family of densities $\rho_{a}$, parameterized by $a\in\mathbb{R}^{N_{a}}$. When employing gradient descent, taking the continuous time limit, the parameter $a$ evolves according to the gradient flow:
(5) $\frac{\mathrm{d}a}{\mathrm{d}t}=-\nabla_{a}{\rm KL}[\rho_{a}\|\rho_{\rm post}].$
The steepest descent direction can be interpreted as
(6) $-\nabla_{a}{\rm KL}[\rho_{a}\|\rho_{\rm post}]=\operatorname*{arg\,min}_{\sigma}\frac{\langle\nabla_{a}{\rm KL}[\rho_{a}\|\rho_{\rm post}],\sigma\rangle}{\sqrt{\langle\sigma,\sigma\rangle}},$
[…]
(7) $-\mathfrak{M}(a)^{-1}\nabla_{a}{\rm KL}[\rho_{a}\|\rho_{\rm post}]=\operatorname*{arg\,min}_{\sigma}\frac{\langle\nabla_{a}{\rm KL}[\rho_{a}\|\rho_{\rm post}],\sigma\rangle}{\sqrt{\langle\sigma,\mathfrak{M}(a)\sigma\rangle}}.$
This modification leads to a different gradient flow for updating $a$ as
(8) $\frac{\mathrm{d}a}{\mathrm{d}t}=-\mathfrak{M}(a)^{-1}\nabla_{a}{\rm KL}[\rho_{a}\|\rho_{\rm post}].$
The present work focuses on the natural gradient [1], where the metric tensor is the Fisher information matrix [53]
$\mathfrak{M}(a)={\rm FIM}(a):=\int\frac{\partial\log\rho_{a}(\theta)}{\partial a}\frac{\partial\log\rho_{a}(\theta)}{\partial a}^{T}\rho_{a}(\theta)\mathrm{d}\theta.$
The Fisher-information matrix is related to the Hessian matrix of the KL-divergence [46], because the Taylor expansion of the KL-divergence (4) between $\rho_{a}$ and its neighbor $\rho_{a+\mathrm{d}a}$, gives
(9) ${\rm KL}[\rho_{a+\mathrm{d}a}\|\rho_{a}]=\frac{1}{2}\mathrm{d}a^{T}{\rm FIM}(a)\mathrm{d}a+o(\lVert da\rVert^{2}).$
Here we used the fact that ${\rm KL}[\rho_{a}\|\rho_{a}]=0$ and $\nabla_{a}{\rm KL}[\rho_{a}\|\rho_{a}]=0$ (the gradient is with respect to the first $\rho_{a}$). Preconditioning the gradient with the Fisher information matrix inherently incorporates geometric information. Consequently, akin to Newton’s method, the Fisher information matrix finds extensive application as a preconditioner to accelerate the optimization process in variational inference. This gives rise to the concept of natural gradient or natural gradient variational inference [1, 41], which corresponds to the following gradient flow:
(10) $\frac{\mathrm{d}a}{\mathrm{d}t}=-{\rm FIM}(a)^{-1}\nabla_{a}{\rm KL}[\rho_{a}\|\rho_{\rm post}].$
Therefore, once the variational family is specified, discretizing the gradient flow introduced above enables the derivation of various practical sampling methods. In what follows, we provide two concrete examples.
2.1 Gaussian Approximation
Gaussian variational inference operates within a Gaussian parametric space, where the variational family
$\rho_{a}^{\rm G}(\theta)=\mathcal{N}(\theta;m,C)$
represents a Gaussian parameterized by its mean $m\in\mathbb{R}^{N_{\theta}}$ and covariance $C\in\mathbb{R}^{N_{\theta}\times N_{\theta}}$, collectively denoted by the parameter vector $a:=[m,C]$. The Fisher information matrix of Gaussian is
(11) ${\rm FIM}(a)=\begin{bmatrix}C^{-1}&\\ &X\end{bmatrix},$
where $X$ is a 4-th order tensor. Its action on any matrix $Y\in\mathbb{R}^{N_{\theta}\times N_{\theta}}$ is given by
(12) $XY=\frac{1}{4}C^{-1}(Y+Y^{T})C^{-1}.$
Incorporating the Fisher information matrix into Eq. 10 and applying Eq. 2 yield the following natural gradient flow
(13) $\dot{m}_{t}=-C_{t}\mathbb{E}_{\rho_{a_{t}}}[\nabla\Phi_{R}],\qquad\dot{C}_{t}=C_{t}-C_{t}\mathbb{E}_{\rho_{a_{t}}}[\nabla^{2}\Phi_{R}]C_{t}.$
By using the fact $\dot{C}_{t}^{-1}=-C_{t}^{-1}\frac{\mathrm{d}C_{t}}{\mathrm{d}t}C_{t}^{-1},$ we can rewrite the covariance evolution equation as
(14) $\dot{C}_{t}^{-1}$ $=\mathbb{E}_{\rho_{a_{t}}}[\nabla^{2}\Phi_{R}]-C_{t}^{-1}.$
The above dynamics Eq. 13 is affine invariant [13, Section 5.4.1]. Consequently, when the posterior is Gaussian, it converges exponentially fast to the posterior at a rate of $\mathcal{O}(e^{-t})$ [13, Theorem 5.6][22, 10, 8], where the exponent of the convergence rate is independent of the posterior. Furthermore, the time discretization of the natural gradient flow Eq. 13 exhibits superior stability. This stability arises from the fact that when the posterior density is log-concave, i.e., when $\nabla^{2}\Phi_{R}$ is positive semidefinite, the forward Euler discretization of Eq. 14 with $10<\Delta t<$ 1 ensures covariance positivity:
$C_{t+\Delta t}^{-1}=C_{t}^{-1}+\Delta t\bigl{(}\mathbb{E}_{\rho_{a_{t}}}[\nabla^{2}\Phi_{R}]-C_{t}^{-1}\bigr{)}=(1-\Delta t)C_{t}^{-1}+\Delta t\mathbb{E}_{\rho_{a_{t}}}[\nabla^{2}\Phi_{R}]\succ 0,$
provided that the approximation of $\mathbb{E}_{\rho_{a_{t}}}[\nabla^{2}\Phi_{R}]$ remains positive semidefinite.
2.2 Gaussian Mixture Approximation
Gaussian mixture variational inference considers the Gaussian mixture parametric space, where the variational family
$\rho_{a}^{\rm GM}(\theta)=\sum_{k=1}^{K}w_{k}\mathcal{N}(\theta;m_{k},C_{k})$
is a $K$-component Gaussian mixture, parameterized by means $m_{k}\in\mathbb{R}^{N_{\theta}}$, covariances $C_{k}\in\mathbb{R}^{N_{\theta}\times N_{\theta}}$ and weights $w_{k}\in\mathbb{R}_{\geq 0}$, collectively denoted by the parameter vector
$a:=[m_{1},...,m_{k},...,m_{K},C_{1}....,C_{k},...,C_{K},w_{1},...,w_{k},...,w_{K}].$
Weights satisfy $\sum_{k=1}^{K}w_{k}=1$. To compute the gradient flow in Eq. 10, we first evaluate the derivatives of the KL divergence in Eq. 4 with respect to $a$:
(15a) $\frac{\partial{\rm KL}[\rho_{a}^{\rm GM}\|\rho_{\rm post}]}{\partial m_{k}}=w_{k}\int\mathcal{N}_{k}(\theta)\Bigl{(}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\Phi_{R}\Bigr{)}\mathrm{d}\theta,$
(15b) $\frac{\partial{\rm KL}[\rho_{a}^{\rm GM}\|\rho_{\rm post}]}{\partial C_{k}}=\frac{w_{k}}{2}\int\mathcal{N}_{k}(\theta)\Bigl{(}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\Bigr{)}\mathrm{d}\theta,$
(15c) $\frac{\partial{\rm KL}[\rho_{a}^{\rm GM}\|\rho_{\rm post}]}{\partial w_{k}}=\int\mathcal{N}_{k}(\theta)\Bigl{(}\log\frac{\rho_{a}^{\rm GM}}{\rho_{\rm post}}+1\Bigr{)}\mathrm{d}\theta.$
Here, we simplify the notation by denoting $\mathcal{N}_{k}(\theta)$ as $\mathcal{N}(\theta;m_{k},C_{k})$. The steepest descent direction is determined by the following constrained optimization problem:
(16) $\operatorname*{arg\,min}_{\sigma}\frac{\langle\nabla_{a}{\rm KL}[\rho_{a}\|\rho_{\rm post}],\sigma\rangle}{\sqrt{\langle\sigma,\textrm{FIM}(a)\sigma\rangle}}\quad\textrm{ s.t. }\quad\sum_{k=1}^{K}\sigma_{\dot{w}_{k}}=0,$
where $\{\sigma_{\dot{w}_{k}}\}$ represents the descent directions corresponding to the weights. The Karush–Kuhn–Tucker conditions of Eq. 16 lead to the following natural gradient flow:
(17) $\begin{bmatrix}\dot{m}_{k}\\ \dot{C}_{k}\\ \dot{w}_{k}\end{bmatrix}=-({\rm FIM}(a))^{-1}\begin{bmatrix}w_{k}\int\mathcal{N}_{k}(\theta)\Bigl{(}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\Phi_{R}\Bigr{)}\mathrm{d}\theta\\ \frac{w_{k}}{2}\int\mathcal{N}_{k}(\theta)\Bigl{(}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\Bigr{)}\mathrm{d}\theta\\ \int\mathcal{N}_{k}(\theta)\bigl{(}\log\rho_{a}^{\rm GM}+\Phi_{R}\bigr{)}\mathrm{d}\theta+\lambda\end{bmatrix}.$
Here $\lambda\in\mathbb{R}$ is the Lagrangian multiplier, determined by the constraint $\sum\dot{w}_{k}=1$. Its value depends on ${\rm FIM}(a)$.
The Fisher information matrix ${\rm FIM}(a)$ for Gaussian mixtures does not have a closed-form expression, and its inversion is computationally expensive. To improve efficiency, diagonal approximations of the Fisher information matrix have been used in the literature [14, Appendix C.8][41], leading to the following approximation:
(18) ${\rm FIM}(a)\approx\textrm{Diag}\left(w_{1}C_{1}^{-1},...,w_{K}C_{K}^{-1},w_{1}X_{1},...,w_{K}X_{K},\frac{1}{w_{1}},...,\frac{1}{w_{K}}\right).$
Here each $X_{k}$ is a 4-th order tensor, and its action on any matrix $Y\in\mathbb{R}^{N_{\theta}\times N_{\theta}}$ is given by
(19) $X_{k}Y=\frac{1}{4}C_{k}^{-1}(Y+Y^{T})C_{k}^{-1}.$
Substituting the approximated Fisher information matrix Eq. 18 into the natural gradient flow Eq. 17 leads to the following equations:
(20) $\begin{split}\dot{m}_{k}&=-C_{k}\int\mathcal{N}_{k}(\theta)\Bigl{(}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\Phi_{R}\Bigr{)}\mathrm{d}\theta,\\ \dot{C}_{k}&=-C_{k}\Bigl{(}\int\mathcal{N}_{k}(\theta)\bigl{(}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{)}\mathrm{d}\theta\Bigr{)}C_{k},\\ \dot{w}_{k}&=-w_{k}\int\Bigl{(}\mathcal{N}_{k}(\theta)-\rho_{a}^{\rm GM}\Bigr{)}\bigl{(}\log\rho_{a}^{\rm GM}+\Phi_{R}\bigr{)}\mathrm{d}\theta.\end{split}$
Here $\lambda=-\int\rho_{a}^{\rm GM}\bigl{(}\log\rho_{a}^{\rm GM}+\Phi_{R}\bigr{)}\mathrm{d}\theta$. Similar to Eq. 14, the covariance evolution equation can be rewritten as:
(21) $\frac{\mathrm{d}C_{k}^{-1}}{\mathrm{d}t}$ $=\int\mathcal{N}_{k}(\theta)\bigl{(}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{)}\mathrm{d}\theta.$
And the weight evolution equation can be reformulated as
(22) $\frac{\mathrm{d}\log w_{k}}{\mathrm{d}t}$ $=-\int\Bigl{(}\mathcal{N}_{k}(\theta)-\rho_{a}^{\rm GM}\Bigr{)}\bigl{(}\log\rho_{a}^{\rm GM}+\Phi_{R}\bigr{)}\mathrm{d}\theta$
to ensure the weights remain positive. In the following, we also refer to Eq. 20 as the natural gradient flow, although involving an approximation of the Fisher information matrix. It is worth noting that when the mode number is $K=1$, the natural gradient flow with Gaussian mixture approximation Eq. 20 reduces to the natural gradient flow with Gaussian approximation Eq. 13.
Designing effective schemes to discretize Eq. 20 remains challenging, particularly due to issues such as the computation of the Hessian matrix, the collapse of different modes, and the singularity of the covariance. A stable, derivative free approximation of Eq. 20 is the main focus of the present work.
3 Derivative Free Gaussian Mixture Variational Inference
In this section, we introduce a stable, derivative free approximation of Eq. 20 for Bayesian inverse problems, termed Derivative Free Gaussian Mixture Variational Inference (DF-GMVI). A key component of practical sampling algorithms is the design of effective quadrature rules. Commonly used methods include the mean-point approximation, unscented transformations [38, 36], cubature transformations [2], the Monte Carlo approach, and stochastic approximation techniqeus [29]. The DF-GMVI method employs two quadrature rules, as defined in Definitions 3.1 and 3.2, based the following considerations:
  1. 1.
    Numerical experiments presented in Appendix A demonstrate that, in this context, the mean-point approximation achieves superior convergence.
  2. 2.
    Building on the mean-point approximation, additional corrections are developed to approximate Hessian expectations, capturing as much curvature information of $\mathcal{F}$ as possible, while maintaining linear complexity in the evaluation of $\mathcal{F}$. These corrections ensure covariance positivity and affine invariance, significantly improving the stability of the overall algorithm, as detailed in Section 4.
We begin by deriving a specialized quadrature rule to compute expectations of the following forms:
(23) $\mathbb{E}_{\mathcal{N}}\bigl{[}\Phi_{R}\bigr{]},\,\mathbb{E}_{\mathcal{N}}\bigl{[}\nabla_{\theta}\Phi_{R}\bigr{]},\,\textrm{and}\,\mathbb{E}_{\mathcal{N}}\bigl{[}\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{]}$
with respect to the Gaussian density $\mathcal{N}(\theta;m,C)$ in a derivative free manner. Here, we assume access only to $\mathcal{F}(\theta)$, and recall that $\Phi_{R}(\theta)=\frac{1}{2}\mathcal{F}(\theta)^{T}\mathcal{F}(\theta)$ has a nonlinear least-squares structure.
Definition 3.1.
Given $\theta\sim\mathcal{N}(\theta;m,C)\in\mathbb{R}^{N_{\theta}}$ and a hyperparameter $\alpha>0$, we generate $12N_{\theta}+$ 1 quadrature points
$\theta_{0}=m\qquad\theta_{i}=m+\alpha[\sqrt{C}]_{i}\qquad\theta_{N_{\theta}+i}=m-\alpha[\sqrt{C}]_{i}\quad(1\leq i\leq N_{\theta}),$
where $\sqrt{C}$ is the square root matrix of $C$, such that $C=\sqrt{C}\sqrt{C}^{T}$. And $[\sqrt{C}]_{i}$ denotes its $i$-th column. Given $\mathcal{F}:\mathbb{R}^{N_{\theta}}\rightarrow\mathbb{R}^{N_{x}}$, we compute vectors
(24) $c=\mathcal{F}(\theta_{0}),\ b_{i}=\frac{\mathcal{F}(\theta_{i})-\mathcal{F}(\theta_{N_{\theta}+i})}{2\alpha},\ a_{i}=\frac{\mathcal{F}(\theta_{i})+\mathcal{F}(\theta_{N_{\theta}+i})-2\mathcal{F}(\theta_{0})}{2\alpha^{2}}\ (1\leq i\leq N_{\theta}),$
and denote $B=[b_{1};\,b_{2};\,\cdots b_{N_{\theta}}]\in\mathbb{R}^{N_{x}\times N_{\theta}}$ and $A=[a_{1};\,a_{2};\,\cdots a_{N_{\theta}}]\in\mathbb{R}^{N_{x}\times N_{\theta}}$. The expectation of the function is then approximated as
(25) $\mathbb{E}_{\mathcal{N}}[\Phi_{R}]=\frac{1}{2}\mathbb{E}_{\mathcal{N}}\Bigl{[}\mathcal{F}(\theta)^{T}\mathcal{F}(\theta)\Bigr{]}\approx\frac{1}{2}c^{T}c.$
The expectation of the gradient is approximated as
(26) $\mathbb{E}_{\mathcal{N}}[\nabla_{\theta}\Phi_{R}]=\frac{1}{2}\mathbb{E}_{\mathcal{N}}\Bigl{[}\nabla_{\theta}\bigl{(}\mathcal{F}(\theta)^{T}\mathcal{F}(\theta)\bigr{)}\Bigr{]}\approx\sqrt{C}^{-T}B^{T}c.$
The expectation of the Hessian is approximated as
(27) $\mathbb{E}_{\mathcal{N}}[\nabla_{\theta}\nabla_{\theta}\Phi_{R}]=\frac{1}{2}\mathbb{E}_{\mathcal{N}}\Bigl{[}\nabla_{\theta}\nabla_{\theta}\bigl{(}\mathcal{F}(\theta)^{T}\mathcal{F}(\theta)\bigl{)}\Bigr{]}\approx\sqrt{C}^{-T}(6\mathrm{Diag}(A^{T}A)+B^{T}B)\sqrt{C}^{-1},$
where $\mathrm{Diag}(\cdot)$ extracts the diagonal elements of a matrix to form a diagonal matrix.
To reach Definition 3.1, we first define ~$\widetilde{\mathcal{F}}(\widetilde{\theta})=\mathcal{F}(m+\sqrt{C}\widetilde{\theta})$. The expectations of $\Phi_{R}(\theta)$ and its gradient in Eqs. 25 and 26 are then approximated using the mean-point approximation with the finite difference approximation of the gradient $\nabla_{\widetilde{\theta}}\widetilde{\mathcal{F}}(0)^{T}$, as given in Eq. 24. The expectation of the Hessian in Eq. 27 can be interpreted as a correction that incorporates curvature information into the mean-point approximation in the Gaussian Newton formulation, $\sqrt{C}^{-T}(B^{T}B)\sqrt{C}^{-1}$. The derivation of the additional terms $\sqrt{C}^{-T}(6\mathrm{Diag}(A^{T}A))\sqrt{C}^{-1}$ proceeds as follows: We assume that ~$\widetilde{\mathcal{F}}(\widetilde{\theta})=A\widetilde{\theta}\odot\widetilde{\theta}+B\widetilde{\theta}+c$, where $\theta\odot\theta$ denoting elementwise product. The column vectors of matrices $A$ and $B$ are approximated by finite differences given in Eq. 24. Under this assumption, the expectation of the Hessian has an analytical form:
$\begin{split}\frac{1}{2}\mathbb{E}_{\mathcal{N}}\Bigl{[}\nabla_{\theta}\nabla_{\theta}\bigl{(}\mathcal{F}(\theta)^{T}\mathcal{F}(\theta)\bigl{)}\Bigr{]}&=\sqrt{C}^{-T}\mathbb{E}_{\widetilde{\mathcal{N}}(0,I)}[\nabla_{\widetilde{\theta}}\nabla_{\widetilde{\theta}}\widetilde{\mathcal{F}}(\widetilde{\theta})^{T}\widetilde{\mathcal{F}}(\widetilde{\theta})+\nabla_{\widetilde{\theta}}\widetilde{\mathcal{F}}(\widetilde{\theta})^{T}\nabla_{\widetilde{\theta}}\widetilde{\mathcal{F}}(\widetilde{\theta})]\sqrt{C}^{-1}\\ &=\sqrt{C}^{-T}(D+2\mathrm{Diag}(A^{T}c)+B^{T}B)\sqrt{C}^{-1},\end{split}$
where $D$ is a diagonal matrix with $D_{jj}=2\sum_{k}a_{j}^{T}a_{k}+4a_{j}^{T}a_{j}$. To ensure covariance positivity, we retain only the positive definite part of this expectation, resulting in Eq. 27.
Moreover, it is worth noticing the quadrature rule approximation Eq. 27 achieves exactness when $\mathcal{F}$ is linear. However, black-box variational inference [52] formulates this term as follows:
$\mathbb{E}_{\mathcal{N}}[\nabla_{\theta}\nabla_{\theta}\Phi_{R}(\theta)]=2\int\nabla_{C}\mathcal{N}(\theta;m,C)\Phi_{R}(\theta)\mathrm{d}\theta=C^{-1}\mathbb{E}_{\mathcal{N}}\Bigl{[}\bigl{(}(\theta-m)(\theta-m)^{T}-C\bigr{)}\Phi_{R}\Bigr{]}C^{-1},$
which is derived using integration by parts. Evaluating this requires computing the expectation of a highly nonlinear matrix function. Even when $\mathcal{F}$ is linear, using either the unscented [38, 36] or cubature [2] transformations does not provide an accurate approximation. And stochastic approximation suffers from high variance, which slows convergence (see Section 5.2).
Then, we develop the quadrature rule for approximating the expectation of $\log\rho^{\rm GM}$ and its derivatives used in Eq. 20.
Definition 3.2.
Given Gaussian mixture $\rho_{a}^{\rm GM}(\theta)=\sum_{k=1}^{K}w_{k}\mathcal{N}(\theta;m_{k},C_{k}),$ the expectation of $\log\rho_{a}^{\rm GM}(\theta)$ and its gradient with respect to its Gaussian component $\mathcal{N}_{k}(\theta)=\mathcal{N}(\theta;m_{k},C_{k})$ are approximated as
(28) $\mathbb{E}_{\mathcal{N}_{k}}[\log\rho_{a}^{\rm GM}(\theta)]\approx\log\rho_{a}^{\rm GM}(m_{k})\qquad\mathbb{E}_{\mathcal{N}_{k}}[\nabla_{\theta}\log\rho_{a}^{\rm GM}(\theta)]\approx\nabla_{\theta}\log\rho_{a}^{\rm GM}(m_{k}).$
The expectation of the Hessian of $\log\rho_{a}^{\rm GM}(\theta)$ with respect to its Gaussian component $\mathcal{N}_{k}$ is approximated as
(29) $\begin{split}&\mathbb{E}_{\mathcal{N}_{k}}\Bigl{[}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(\theta)\Bigr{]}\\ &=-\mathbb{E}_{\mathcal{N}_{k}}\Bigl{[}\frac{\Bigl{(}\sum_{i}w_{i}v_{i}\mathcal{N}_{i}(\theta)\Bigr{)}\Bigl{(}\sum_{i}w_{i}v_{i}\mathcal{N}_{i}(\theta)\Bigr{)}^{T}}{\rho_{a}^{\rm GM}(\theta)^{2}}+\frac{\sum_{i}w_{i}\Bigl{(}C_{i}^{-1}-v_{i}v_{i}^{T}\Bigr{)}\mathcal{N}_{i}(\theta)}{\rho_{a}^{\rm GM}(\theta)}\Bigr{]}\\ &=\mathbb{E}_{\mathcal{N}_{k}}\Bigl{[}\frac{\sum_{i<j}w_{i}w_{j}\bigl{(}v_{i}-v_{j}\bigr{)}\bigl{(}v_{i}-v_{j}\bigr{)}^{T}\mathcal{N}_{i}(\theta)\mathcal{N}_{j}(\theta)}{\rho_{a}^{\rm GM}(\theta)^{2}}\Bigr{]}-\mathbb{E}_{\mathcal{N}_{k}}\Bigl{[}\frac{\sum_{i}w_{i}\mathcal{N}_{i}(\theta)C_{i}^{-1}}{\rho_{a}^{\rm GM}(\theta)}\Bigr{]}\\ &\approx\frac{\sum_{i<j}w_{i}w_{j}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}^{T}\mathcal{N}_{i}(m_{k})\mathcal{N}_{j}(m_{k})}{\rho_{a}^{\rm GM}(m_{k})^{2}}-C_{k}^{-1}.\end{split}$
Here we denote $v_{i}(\theta)=C_{i}^{-1}(\theta-m_{i})$.
The proposed quadrature rule in Definition 3.2 is consistent with Definition 3.1, which is important because the variational inference relies on the balance of these two terms. Specifically, the expectation of $\log\rho_{a}^{\rm GM}(\theta)$ and its gradient are computed using the mean-point approximation. For the Hessian approximation in Eq. 29, it is decomposed into two terms. The first term is approximated using the mean-point approximation. The second term is approximated as $C_{k}^{-1}$, which is a correction to the mean-point approximation. This correction satisfies the following equality:
$\sum_{k}w_{k}\mathbb{E}_{\mathcal{N}_{k}}\Bigl{[}\frac{\sum_{i}w_{i}\mathcal{N}_{i}(\theta)C_{i}^{-1}}{\rho_{a}^{\rm GM}(\theta)}\Bigr{]}=\int\sum_{i}w_{i}\mathcal{N}_{i}(\theta)C_{i}^{-1}\mathrm{d}\theta=\sum_{k}w_{k}C_{k}^{-1}.$
When $K=1$, this correction becomes exact. This correction ensures covariance positivity, significantly enhancing the stability of the algorithm (see Proposition 4.1).
Finally, we update the covariances, means, and weights sequentially using a forward Euler scheme from time $t$ to time $t+\Delta t$ with time step $\Delta t$. For the mean update, we use the updated covariance $C_{k}(t+\Delta t)$ in a Gauss–Seidel manner on the right-hand side.
(30) $\begin{split}C_{k}^{-1}(t+\Delta t)&=C_{k}^{-1}(t)+\Delta t~{}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{\}},\\ m_{k}(t+\Delta t)&=m_{k}(t)-\Delta t~{}C_{k}(t+\Delta t)~{}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\Phi_{R}\bigr{\}},\\ \log w_{k}(t+\Delta t)&=\log w_{k}(t)-\Delta t~{}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\log\rho_{a}^{\rm GM}(t)+\Phi_{R}\bigr{\}}.\end{split}$
Here, $\mathrm{\mathsf{QR}}$ denotes the quadrature rules as defined in Definitions 3.1 and 3.2, and $\mathcal{N}_{k}$ denotes the $k$-th Gaussian component at time $t$. We then normalize ${w_{k}(t+\Delta t)}_{k=1}^{K}$ and, for efficiency, set a lower bound of $w_{k}$ at a default value of $10^{-8}$ during normalization. The computational cost with respect to evaluating $\mathcal{F}$ is $(2N_{\theta}+1)KN_{t}$. This is because evaluating the quadrature rule in Definition 3.1 requires computing $\mathcal{F}$ at $12N_{\theta}+$ 1 quadrature points for each Gaussian mode, where $N_{t}$ represents the total number of time steps. It is worth mentioning the evaluation of $\mathcal{F}$ at the $(2N_{\theta}+1)K$ quadrature points can be performed in an embarrassingly parallel manner.
4 Theoretical Insight
In this section, we present the theoretical insights underlying the proposed DF-GMVI method, including its covariance positivity-preserving and affine invariance properties. These insights are also applied to design adaptive time-stepping schemes and select gradient flows, which significantly enhance the performance of other Gaussian mixture variational inference methods (see Section 5.2).
A common issue with Gaussian mixture-based sampling methods is instability due to the covariance matrix becoming singular or non-positive. The following proposition establishes a condition on the time step $\Delta t$ that ensures the covariance matrix remains positive definite during the update process, which relies on our design of the quadrature rules used for approximating Hessian expectations in Section 3. The proposition guarantees that we can choose a sufficiently large $\Delta t$, enabling faster convergence without introducing numerical instability (see Section 5).
Proposition 4.1.
For the DF-GMVI algorithm described in Eq. 30, if $10<\Delta t<$ 1, then $C_{k}$ remains positive definite.
Proof 4.2.
Incorporating the quadrature rules from Eq. 27 and Eq. 29 into Eq. 30, the update rule for $C_{k}$ becomes
(31) $\begin{split}C_{k}^{-1}(t+\Delta t)=&C_{k}^{-1}(t)+\Delta t~{}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{\}}\\ =&C_{k}^{-1}(t)+\Delta t~{}\Bigl{(}\sqrt{C}^{-T}(6\mathrm{Diag}(A^{T}A)+B^{T}B)\sqrt{C}^{-1}-C_{k}^{-1}\\ &+\frac{\sum_{i<j}w_{i}w_{j}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}^{T}\mathcal{N}_{i}(m_{k})\mathcal{N}_{j}(m_{k})}{\rho_{a}^{\rm GM}(m_{k})^{2}}\Bigr{)}\\ =&(1-\Delta t)C_{k}^{-1}(t)+\Delta t~{}\Bigl{(}\sqrt{C}^{-T}(6\mathrm{Diag}(A^{T}A)+B^{T}B)\sqrt{C}^{-1}\\ &+\frac{\sum_{i<j}w_{i}w_{j}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}\bigl{(}v_{i}(m_{k})-v_{j}(m_{k})\bigr{)}^{T}\mathcal{N}_{i}(m_{k})\mathcal{N}_{j}(m_{k})}{\rho_{a}^{\rm GM}(m_{k})^{2}}\Bigr{)}.\end{split}$
Since the matrix within the brackets is always positive semidefinite and $(1-\Delta t)C_{k}^{-1}(t)$ is positive definite, we conclude that $C_{k}^{-1}(t+\Delta t)$ is also positive definite.
Moreover, sampling methods are said to be affine invariant if their behavior remains unchanged under any invertible affine transformation. This property makes these methods especially effective for highly anisotropic posterior distributions [27, 21]. Their effectiveness arises from consistent performance across coordinate systems related by affine transformations. Specifically, the convergence properties of these methods can be understood by examining the optimal coordinate system, which minimizes anisotropy to the fullest extent, across all affine transformations. The following proposition establishes a restricted affine invariance of DF-GMVI method.
Proposition 4.3.
The sampling algorithm applied to the posterior distribution, which is proportional to $\mathrm{e}^{-\Phi_{R}(\theta)}$, produces Gaussian mixture approximations parameterized by $\{m_{k}(t),C_{k}(t),w_{k}(t)\}_{k=1}^{K}$. When the algorithm is applied to the transformed posterior, with $\widetilde{\Phi}_{R}(\widetilde{\theta})=\Phi_{R}(T^{-1}(\widetilde{\theta}-d))$, under any invertible affine mapping $\varphi:\theta\rightarrow\widetilde{\theta}=T\theta+d$, where $T\in\mathcal{T}$ is an invertible matrix and $d$ is a translation vector, it produces updated parameters $\{\widetilde{m}_{k}(t),\widetilde{C}_{k}(t),\widetilde{w}_{k}(t)\}_{k=1}^{K}$. The algorithm is said to be $\mathcal{T}$-invariant if the following holds:
$\widetilde{w}_{k}(t)=w_{k}(t),\quad\widetilde{m}_{k}(t)=Tm_{k}(t)+d,\quad\widetilde{C}_{k}(t)=TC_{k}(t)T^{T}.$
Then, we have:
  1. 1.
    The natural gradient flow defined by Eq. 20 is affine invariant, i.e., $\mathcal{T}$ contains all invertible matrices.
  2. 2.
    The DF-GMVI algorithm Eq. 30 is $\mathcal{T}$-invariant, when using Cholesky decomposition to compute the square root matrix $\sqrt{C}$ in Definition 3.1, where $\mathcal{T}$ denotes the group of invertible lower triangular matrices.
We provide a proof of this proposition in Appendix B. The affine invariance is restricted because the Cholesky decomposition of the covariance is only invariant for lower triangular matrix group. The affine invariance enables the DF-GMVI method and variational inference methods based on natural gradient flow (20) to converge faster than other gradient flows with Gaussian mixture approximations, as further demonstrated numerically in Section 5.2.
5 Numerical Study
In this section, we present numerical studies regarding the proposed DF-GMVI method. We focus on posterior distributions of unknown parameters or fields arising in inverse problems that may exhibit multiple modes. Three types of model problems are considered:
  1. 1.
    A one-dimensional bimodal problem from [14]: we use this problem as a proof-of-concept example. Our result shows that the convergence rate remains unchanged no matter how overlapped the two modes are. Additionally, DF-GMVI outperforms the derivative free Gaussian mixture Kalman inversion presented in [14].
  2. 2.
    Several multi-dimensional problems: we use these problems to demonstrate that DF-GMVI is robust with respect to the posterior featuring multiple modes, infinite many modes, and modes with narrow and curved shapes. And the convergence rate remains robust across different problem dimensionalities. Additionally, we compare DF-GMVI with other state-of-the-art sampling approaches, including Gaussian mixture variational inference methods leveraging gradient or Hessian information, BBVI, and MCMC methods, to highlight its strengths. Notably, the theoretical insights presented in Section 4 are general and are used to design adaptive time-stepping schemes that significantly enhance the robustness of these comparable Gaussian mixture variational inference methods.
  3. 3.
    Navier-Stokes problem: we consider the inverse problem of recovering the initial velocity field of a Navier-Stokes flow. The problem is structured to exhibit symmetry, resulting in two modes in the posterior distribution. We demonstrate that DF-GMVI effectively captures both modes, highlighting its potential for tackling multimodal problems in large-scale, high-dimensional applications.
For all tests, we use $\Delta t=0.5$ and run 200 iterations, resulting in a total time of $T=100$. We set $\alpha=10^{-3}$ in the derivative free quadrature rule in Definition 3.1. Our code is available online: https://github.com/PKU-CMEGroup/InverseProblems.jl/tree/master/Derivative-Free-Variational-Inference.
5.1 One-Dimensional Problems
In this subsection, we consider the 1D bimodal inverse problem from [14], associated with the forward model
$y=\mathcal{G}(\theta)+\eta\quad\textrm{ with }\quad y=1,\quad\mathcal{G}(\theta)=\theta^{2}.$
The prior is $\rho_{\rm prior}\sim\mathcal{N}(3,2^{2})$ and consider different noise levels:
[…]
Note that the overlap between these two modes increases as the noise level rises. In case A, the two modes are well separated, while in case D, they are nearly indistinguishable (See Fig. 1).
[…]
5.2 Multi-Dimensional Problems
In this subsection, we first examine several 2D sampling problems, along with their 100-dimensional modified versions. We then conclude with a comprehensive comparison to other state-of-the-art sampling approaches. For the 2D sampling problems, we define $\Phi_{R}(\theta)=\frac{1}{2}\mathcal{F}(\theta)^{T}\mathcal{F}(\theta)$, where $\theta=[\theta_{(1)},\theta_{(2)}]^{T}\in\mathbb{R}^{2}$, and $\mathcal{F}(\theta)$ is defined as follows.
  1. Case A
    : The distribution is a Gaussian with
    $\mathcal{F}(\theta)=y-A\theta\quad A=\begin{bmatrix}1&1\\ 1&2\end{bmatrix}\quad\textrm{ and }\quad y=\begin{bmatrix}0\\ 1\end{bmatrix}.$
  2. Case B
    : The distribution has four modes with different weights [14, Appendix D]
    $\quad\mathcal{F}(\theta)=y-\begin{bmatrix}(\theta_{(1)}-\theta_{(2)})^{2}\\ (\theta_{(1)}+\theta_{(2)})^{2}\\ \theta_{(1)}\\ \theta_{(2)}\end{bmatrix}\quad\textrm{ and }\quad y=\begin{bmatrix}4.2297\\ 4.2297\\ 0.5\\ 0.0\end{bmatrix}.$
  3. Case C
    : The distribution is circular in shape [14, Appendix E], with its mass concentrated along an elliptical region,
    $\mathcal{F}(\theta)=2(y-\theta^{T}A\theta)\quad\textrm{ and }\quad A=\begin{bmatrix}1&1\\ 1&2\end{bmatrix}\quad y=\begin{bmatrix}1\end{bmatrix}.$
  4. Case D
    : The distribution follows the Rosenbrock function, which has a characteristic “banana” shape [27], with
    $\mathcal{F}(\theta)=\frac{1}{\sqrt{10}}\Bigl{(}y-\begin{bmatrix}10(\theta_{(2)}-\theta_{(1)}^{2})\\ \theta_{(1)}\end{bmatrix}\Bigr{)}\quad\textrm{ and }\quad y=\begin{bmatrix}0\\ 1\end{bmatrix}.$
  5. Case E
    : The distribution is an extension of the Rosenbrock function [15] and is characterized by its bimodal, “banana” shaped density.
    $\mathcal{F}(\theta)=y-\begin{bmatrix}\log\bigl{(}100(\theta_{(2)}-\theta_{(1)}^{2})^{2}+(1-\theta_{(1)})^{2}\bigr{)}/0.3\\ \theta_{(1)}\\ \theta_{(2)}\end{bmatrix}\quad\textrm{ and }\quad y=\begin{bmatrix}\log(101)\\ 0\\ 0\end{bmatrix}.$
We apply DF-GMVI with $K=10,20$, and $40$ modes, each randomly initialized as $\mathcal{N}(0,I)$ with equal weights. The density estimations at the 200th iteration are shown in Fig. 2, with each row corresponding to a case (A to D). In each row, the reference density is displayed in the first panel, followed by the results from DF-GMVI with $K=10,20$, and $40$ modes. The estimated densities closely match the reference. The fourth panel in each row shows the error in terms of total variation over the iterations. DF-GMVI effectively captures multiple modes (Cases B and E), densities characterized by narrow, curved valleys (Cases D and E), which typically pose challenges for sampling methods, as well as densities with infinitely many maximum a posteriori points (Case C). The fourth panel demonstrates that DF-GMVI converges efficiently across all cases. However, it is important to note that when the posterior is Gaussian (Case A), DF-GMVI does not achieve exponential convergence due to the interactions among different Gaussian components; in contrast, Gaussian variational inference Eq. 13[13] achieves exponential convergence in this setting. Furthermore, when the number of modes $K$ is small, or when the initial Gaussian components are distant from the target mode, DF-GMVI may fail to capture all modes of the posterior density. Increasing $K$ helps mitigate this issue and enhances performance.
Figure 2: Multi-dimensional problems with dimensionality $2$, arranged from Case A to Case E from the top row to the bottom row. Each panel shows the reference density alongside the densities estimated by DF-GMVI with $K=10,20$ and $40$ (from left to right). The mean $m_{k}$ of each Gaussian component is marked by red circles, and the initial means are marked by gray crosses. The fourth panel shows the total variation distance between the reference density and the densities estimated by the DF-GMVI over the iterations.
We then modify these sampling problems to a $100$-dimensional problem ($N_{\theta}=100$) by introducing $N_{\theta}-2$ additional variables, $\theta^{c}$. The reference density is defined with
$\Phi_{R}(\theta,\theta^{c})=\frac{1}{2}\mathcal{F}(\theta)^{T}\mathcal{F}(\theta)+\frac{1}{2}(\theta^{c}-K\theta)^{T}(\theta^{c}-K\theta),$
where $\theta\in\mathbb{R}^{2}$, $\theta^{c}\in\mathbb{R}^{N_{\theta}-2}$, and $K\in\mathbb{R}^{(N_{\theta}-2)\times 2}$ is an all-ones matrix. The function $\mathcal{F}$ is defined as in the previous 2-dimensional densities. These high-dimensional densities are constructed so that the marginal densities of $\theta$ remain computable and identical to that of the previous 2-dimensional densities, which serves as the error indicator for evaluating DF-GDVI.
Similarly, we apply DF-GMVI with $K=10,20$, and $40$ modes, each randomly initialized as $\mathcal{N}(0,I)$ with equal weights. The marginal densities obtained from the estimated 100-dimensional densities at the 200th iteration are shown in Fig. 3, using the same configuration as in Fig. 2. The estimated marginal densities closely align with the reference, and the total variation distances are similar to those from the 2-dimensional problems in Fig. 2. This confirms that the performance of DF-GMVI remains robust across different dimensionalities. We conjecture that the difficulty of the sampling problem depends primarily on the complexity of the target density (e.g., the number of modes) rather than the dimensionality.
Figure 3: Multi-dimensional problems with dimensionality 100, arranged from Case A to Case E from the top row to the bottom row. Each panel shows the marginal density of the reference alongside the marginal densities estimated by DF-GMVI with $K=10,20$, and $40$ (from left to right). The projected means of each Gaussian component are marked by red circles, and the projected initial means are marked by grey crosses. The fourth panel displays the total variation between the reference marginal density and the estimated marginal densities across the iterations.
Finally, we compare DF-GMVI with other state-of-the-art methods, including the Gaussian mixture approximation of the natural gradient flow (NGF-VI) Eq. 20 [41], its diagonal covariance variant (NGF-VI-D) [49], the Wasserstein gradient flow (WGF-VI)[40], BBVI [52], and the affine invariant MCMC method [27, 21]. NGF-VI, NGF-VI-D, and WGF-VI are not derivative free, requiring the computation of the gradient or Hessian of $\Phi_{R}$. Expectations are approximated using the mean-point approximation discussed in Appendix A. Motivated by the covariance positivity insight outlined in Section 4, for the NGF-VI and NGF-VI-D, we implement an adaptive time-stepping scheme defined as
$\Delta t=\min\{\Delta t_{\max},\frac{\beta}{\max_{k}\lVert C_{k}\mathbb{E}_{\mathcal{N}_{k}}[\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}]\rVert_{2}}\}$
with $\Delta t_{\max}=0.5$ and $\beta=0.99$, ensuring that the covariance update $C_{k}^{-1}(t+\Delta t)=C_{k}^{-1}(t)+\Delta t\mathbb{E}_{\mathcal{N}_{k}}[\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}]$ maintains positive definiteness. For the WGF-VI, the covariance update equation can be written as
$\dot{C}_{k}^{-1}=C_{k}^{-1}\mathbb{E}_{\mathcal{N}_{k}}+\mathbb{E}_{\mathcal{N}_{k}}C_{k}^{-1},\quad\mathbb{E}_{\mathcal{N}_{k}}=\mathbb{E}_{\mathcal{N}_{k}}[\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(t)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}].$
We follow the update method proposed in [40], $C_{k}^{-1}(t+\Delta t)=M(t)C_{k}^{-1}(t)M(t)$, $M(t)=I+\Delta t\mathbb{E}_{k}(t)$, where $I$ is the identity matrix, to achieve first-order accuracy while ensuring covariance positivity. The time step $\Delta t$ is manually chosen to ensure stability, with values $1.4\times 10^{-1}$, $5.0\times 10^{-3}$, $5.0\times 10^{-3}$, $4.0\times 10^{-3}$ and $8.0\times 10^{-4}$ for Case A to Case E, respectively. For the BBVI, we adopt the natural gradient flow (20), derive the mean and covariance update equations by using integration by parts, as follows:
(32) $\begin{split}\dot{m}_{k}&=-\mathbb{E}_{\mathcal{N}_{k}}[\bigl{(}\theta-m_{k}\bigr{)}\bigl{(}\log\rho_{a}^{\rm GM}+\Phi_{R}\bigr{)}],\\ \dot{C}_{k}&=-\mathbb{E}_{\mathcal{N}_{k}}\Bigl{[}\bigl{(}(\theta-m_{k})(\theta-m_{k})^{T}-C_{k}\bigr{)}(\log\rho_{a}^{\rm GM}+\Phi_{R})\Bigr{]}.\end{split}$
We apply the Monte Carlo method with $5$ ensembles (same as DF-GMVI) to compute these integrals. We use empirical covariance instead of $C_{k}$ in the right-hand-side of the covariance update equation to reduce variance, and implement a similar adaptive time-stepping scheme defined as
$\Delta t=\min\{\Delta t_{\max},\frac{\beta}{\max_{k}\lVert\mathbb{E}_{\mathcal{N}_{k}}\bigl{[}\bigl{(}(\theta-m_{k})(\theta-m_{k})^{T}-C_{k}\bigr{)}(\log\rho_{a}^{\rm GM}+\Phi_{R})\bigr{]}C_{k}^{-1}\rVert_{2}}\}$
with $\Delta t_{\max}=0.5$ and $\beta=0.99$. Finally, for the affine invariant MCMC method, we use the stretch move method proposed in [27]. To recover the equilibrium distributions, we apply Gaussian kernel density estimation with a selectively chosen bandwidth using particles sampled from the last 10 iterations.
Fig. 4 shows the results obtained using Gaussian mixture variational inference approaches with $K=40$ and the affine invariant MCMC method with $J=10^{3}$ particles for the 2-dimensional problems. We observe that these variational inference approaches require very small time steps, particularly in Case E, to ensure stability. Our proposed adaptive time steppers significantly improve this. The BBVI method exhibits random noise, which can be mitigated by increasing the ensemble size for Monte Carlo integrals. Among the variational inference methods, those based on natural gradient flows with affine invariance outperform their counterparts based on Wasserstein gradient flows. Overall, the results obtained by DF-GMVI, as shown in Fig. 2, are similarly comparable and avoid the use of derivatives. In high dimensional spaces, the volume grows exponentially, which makes it difficult for random derivative free methods to explore the space efficiently. As a result, the performance of BBVI and derivative free MCMC methods typically deteriorates with increasing dimensionality, as illustrated in Fig. 5. However, the quadrature rule Definition 3.1 leverages gradient structure, allowing our DF-GMVI method to avoid this issue even in the 100-dimensional problem. Consequently, our DF-GMVI method outperforms the other approaches on the test problems presented in this subsection.
Figure 4: Comparison study for multi-dimensional problems with dimensionality $2$, arranged from Case A to Case E (from top to bottom). Each panel displays the reference density alongside the densities estimated by variational inference methods and the affine invariant MCMC (left to right: NGF-VI, NGF-VI-D, WGF-VI, BBVI, MCMC). The last figure shows the total variation distance between the reference density and the densities estimated by different methods over the iterations.
Figure 5: Comparison study for multi-dimensional problems with dimensionality $100$ (Case B). The figures display the marginal reference density and the estimated marginal densities obtained by BBVI with $K=40$ and ensemble sizes $J=64,256$, and $1024$ for Monte Carlo integration, and by affine invariant MCMC method with ensemble sizes $J=2^{10},2^{12}$, and $2^{14}$ (from left to right). The last figure shows the total variation distance between the marginal reference density and the estimated marginal densities over iterations.
5.3 High-Dimensional Inverse Problem
In this subsection, we study the problem of recovering the initial vorticity field, $\omega_{0}$, of a fluid flow based on measurements taken at later times. The flow is governed by the 2D incompressible Navier-Stokes equation on a periodic domain $D=[0,2\pi]\times[0,2\pi]$, written in the vorticity-streamfunction $\omega-\psi$ formulation:
(33) $\begin{split}&\frac{\partial\omega}{\partial t}+(v\cdot\nabla)\omega-\nu\Delta\omega=\nabla\times f,\\ &\omega=-\Delta\psi\qquad\frac{1}{4\pi^{2}}\int\psi=0\qquad v=[\frac{\partial\psi}{\partial x_{2}},-\frac{\partial\psi}{\partial x_{1}}]^{T}+v_{b}.\end{split}$
Here $v$ denotes the velocity vector, $\nu=0.01$ denotes the viscosity, $v_{b}=[0,2\pi]^{T}$ denotes the non-zero mean background velocity, and $f(x)=[0,\cos(4x_{(1)})]^{T}$ denotes the external forcing.
The problem setup follows [14], which is spatially symmetric with respect to $x_{(1)}=\pi$. The source of the fluid is chosen such that ×$\nabla\times f([x_{(1)},x_{(2)}]^{T})=-\nabla\times f([2\pi-x_{(1)},x_{(2)}]^{T}).$ The observations in the inverse problem are chosen as the difference of pointwise measurements of the vorticity value $\omega([x_{(1)},x_{(2)}]^{T})-\omega([2\pi-x_{(1)},x_{(2)}]^{T})$ at $56$ equidistant points in the left domain (see Figure 6), at $T=0.25$ and $T=0.5$, corrupted with observation error $\eta\sim\mathcal{N}(0,0.1^{2}I)$. Under this set-up, both $\omega_{0}([x_{(1)},x_{(2)}]^{T})$ and $-\omega_{0}([2\pi-x_{(1)},x_{(2)}]^{T})$ will lead to the same measurements. Thus the posterior of the Bayesian inverse problem will be at least bi-modal.
We assume the prior of $\omega_{0}(x,\theta)$ is a Gaussian field with covariance operator $C=(-\Delta)^{-2}$, subject to periodic boundary conditions, on the space of mean zero functions. Following [28], let $\mathbb{Z}_{+}^{2}=\{(l_{1},l_{2})\in\mathbb{Z}^{2},l_{2}>0\}\cup\{(l_{1},0)\in\mathbb{Z}^{2},l_{1}>0\}$ and $\mathbb{Z}_{-}^{2}=-\mathbb{Z}_{+}^{2}$, the corresponding KL expansion of the initial vorticity field is given by
(34) $\omega_{0}(x,\theta)=\sum_{l\in\mathbb{Z}^{2}\backslash\{(0,0)\}}\theta_{(l)}\sqrt{\lambda_{l}}\psi_{l}(x),$
where eigenpairs are of the form
$\psi_{l}(x)=\begin{cases}\frac{\sin(l\cdot x)}{\sqrt{2}\pi}&l\in\mathbb{Z}_{+}^{2}\\ \frac{\cos(l\cdot x)}{\sqrt{2}\pi}&l\in\mathbb{Z}_{-}^{2}\end{cases}\quad\lambda_{l}=\frac{1}{|l|^{4}},$
and $\theta_{(l)}\sim\mathcal{N}(0,2\pi^{2})$. The KL expansion Eq. 34 can be rewritten as a sum over positive integers, where the eigenvalues $\lambda_{l}$ are in descending order. We truncate the expansion to the first $128$ terms and generate the truth vorticity field $\omega_{0}(x;\theta_{\rm ref})$ with $\theta_{\rm ref}\in\mathbb{R}^{128}$; we aim to recover the parameters based on observation data.
Figure 6: The vorticity field $\omega$ at $T=0.25$ and $T=0.5$ and observations $\omega([x_{(1)},x_{(2)}]^{T})-\omega([2\pi-x_{(1)},x_{(2)}]^{T})$ at $56$ equidistant points (solid black dots). Their mirroring points are marked (empty black dots).
We apply DF-GMVI with $K=3$ modes, initialized randomly according to the prior distribution and assigned equal weights. In Fig. 7, we display the true initial vorticity field, $\omega_{0}(x;\theta_{\rm ref})$, along with its mirrored field (where the mirroring of the velocity field induces antisymmetry in the vorticity field), and the three recovered initial vorticity fields, $\omega_{0}(x;m_{k})$, obtained by DF-GMVI at the 200th iteration. Modes 1 and 2 capture $\omega_{0}(x;\theta_{\rm ref})$, while mode 3 captures the mirrored field of $\omega_{0}(x;\theta_{\rm ref})$ itself. Figure 8 illustrates the relative errors in the vorticity field, optimization errors $\Phi_{R}(m_{k})$, the Frobenius norm $\lVert C_{k}\rVert_{F}$, and the Gaussian mixture weights $w_{k}$ (left to right) at each iteration. This demonstrates that DF-GMVI converges in fewer than 50 iterations. Figure 9 shows the marginal distributions of the estimated posterior for the first 16 $\theta_{(l)}$-coefficients obtained by DF-GMVI. The marginal distributions reveal clear bimodality, with the approximate posterior placing high probability mass around the true coefficients while also capturing the alternative possibility (the mirrored value). This example demonstrates DF-GMVI’s potential for effectively handling multimodal posteriors in large-scale, high-dimensional applications.
Figure 7: The true initial vorticity field $\omega_{0}(x;\theta_{\rm ref})$, and recovered initial vorticity fields $\omega_{0}(x;m_{k})$ obtained by DF-GMVI.
Figure 8: Navier-Stokes flow problem: the relative errors of the initial vorticity field, the optimization errors $\Phi_{R}(m_{k})$, the Frobenius norm $\lVert C_{k}\rVert_{F}$, and the Gaussian mixture weights $w_{k}$ (from left to right) for different modes over iterations.
Figure 9: Navier-Stokes flow problem: the true Karhunen-Loeve expansion parameters $\theta_{(i)}$ (black crosses), and mean estimations of $\theta_{(i)}$ for each modes (circles) and the associated marginal distributions obtained by DF-GMVI at the $50$th iteration.
6 Conclusion
In this paper, we have explored guidelines, such as covariance positivity and affine invariance, to design stable derivative free Gaussian variational methods. Based on these guidelines, we have introduced DF-GMVI as an effective approach for solving Bayesian inverse problems. Our numerical experiments demonstrate the superior stability of DF-GMVI and its capability to approximate complex high dimensional posterior distributions. There are several promising directions for future research. On the algorithmic side, developing stable derivative free variational inference methods beyond Bayesian inverse problems Eq. 2 would be valuable. On the theoretical side, a thorough analysis of the stability and convergence properties of the natural gradient flow Eq. 20 with Gaussian mixtures and DF-GMVI for general posterior densities, including log-concave densities, could provide important insights for their practical applications.
Acknowledgments
DZH acknowledges the high-performance computing platform at Peking University. YC acknowledges support from the Courant Instructorship.
Appendix A Quadrature Rule Comparison
In this section, we compare various quadrature rules, including mean-point approximation, unscented [38, 36] and cubature [2] transformations, for approximating Gaussian integrations in solving the natural gradient flow Eq. 20. We examine a 2-dimensional Gaussian mixture target density, with $\Phi_{R}(\theta)=-\log\Bigl{(}\sum_{i=1}^{3}w_{i}\mathcal{N}(\theta;m_{i},C_{i})\Bigr{)},$ where $w_{1}=0.2,\ m_{1}=\begin{bmatrix}1\\ 2\end{bmatrix},\ C_{1}=I$, $w_{2}=0.5,\ m_{2}=\begin{bmatrix}2\\ 1\end{bmatrix},\ C_{2}=I$, and $w_{3}=0.3,\ m_{3}=\begin{bmatrix}-1\\ -1\end{bmatrix},\ C_{3}=I/2$. We implement gradient and Hessian of $\Phi_{R}(\theta)$, and consider four combinations for the quadrature integrations, (1) both $\Phi_{R}$ and $\log\rho_{a}$ related terms integrated with mean-point approximation, (2) both $\Phi_{R}$ and $\log\rho_{a}$ related terms integrated with unscented transformation using weight $\sqrt{3}$ [38, 36], (3) both $\Phi_{R}$ and $\log\rho_{a}$ related terms integrated with the third-order cubature transformation [2], and (4) $\Phi_{R}$ related terms integrated with mean-point approximation and $\log\rho_{a}$ related terms with third-order cubature transformation. We use $K=10$, $\Delta t=0.02$, and solving for 1000 iterations, the convergence results are presented in Fig. 10. In the context of solving gradient flow, the mean-point approximation achieves superior convergence, even though it is not the most accurate quadrature rule. And hence the mean-point approximation is used in Definitions 3.1 and 3.2.
Figure 10: Quadrature rule comparison for the Gaussian mixture target density problem. Each panel displays the reference density and densities estimated by approximating natural gradient flow Eq. 20 with various quadrature rule combinations (from left to right). The mean $m_{k}$ (red circles) and initial mean (grey crosses) of each Gaussian component are marked. The last panel shows the total variation distance between the reference density and the densities estimated.
Appendix B Proof of Proposition 4.3
Proof B.1.
We note that
$\nabla_{\theta}\Phi_{R}(\theta)=T^{T}\nabla_{\widetilde{\theta}}\widetilde{\Phi}_{R}(\widetilde{\theta}),\quad\nabla_{\theta}\nabla_{\theta}\Phi_{R}(\theta)=T^{T}\nabla_{\widetilde{\theta}}\nabla_{\widetilde{\theta}}\widetilde{\Phi}_{R}(\widetilde{\theta})T.$
  1. 1.
    To prove the affine invariance of the natural gradient flow defined by Eq. 20, it suffices to verify that under the affine mapping
    $\widetilde{w}_{k}=w_{k},\quad\widetilde{m}_{k}=Tm_{k}+d,\quad\widetilde{C}_{k}=TC_{k}T^{T},$
    $\widetilde{w}_{k},\widetilde{m}_{k}$ and $\widetilde{C}_{k}$ also satisfy the corresponding equation Eq. 20. We simplify the notation by $\widetilde{\mathcal{N}}_{k}(\widetilde{\theta})=\mathcal{N}(\widetilde{\theta},\widetilde{m_{k}},\widetilde{C_{k}}),\,\rho_{\widetilde{a}}^{\mathrm{GM}}(\widetilde{\theta})=\sum_{k=1}^{K}\widetilde{w}_{k}\widetilde{\mathcal{N}}_{k}(\widetilde{\theta})$. We have $\widetilde{\mathcal{N}}_{k}(\widetilde{\theta})=|\det T|^{-1}\mathcal{N}_{k}(\theta),\,\rho_{\widetilde{a}}^{\mathrm{GM}}(\widetilde{\theta})=|\det T|^{-1}\rho_{a}^{\mathrm{GM}}(\theta)$, and thus $\nabla_{\widetilde{\theta}}\log\rho_{\widetilde{a}}^{\rm GM}(\widetilde{\theta})=T^{-T}\nabla_{\theta}\log\rho_{a}^{\rm GM}(\theta)$, $\nabla_{\widetilde{\theta}}\nabla_{\widetilde{\theta}}\log\rho_{\widetilde{a}}^{\rm GM}(\widetilde{\theta})=T^{-T}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(\theta)T^{-1}$. By these equations and change-of-variable formula, we have that
    $\dot{\widetilde{m}}_{k}=T\dot{m}_{k}$ $=-TC_{k}\int\mathcal{N}_{k}(\theta)\Bigl{(}\nabla_{\theta}\log\rho_{a}^{\rm GM}(\theta)+\nabla_{\theta}\Phi_{R}(\theta)\Bigr{)}\mathrm{d}\theta$
    $=-TC_{k}T^{T}\int\widetilde{\mathcal{N}}_{k}(\widetilde{\theta})T^{-T}\Bigl{(}\nabla_{\theta}\log\rho_{a}^{\rm GM}(\theta)+\nabla_{\theta}\Phi_{R}(\theta)\Bigr{)}\mathrm{d}\widetilde{\theta}$
    $=-\widetilde{C}_{k}\int\widetilde{\mathcal{N}}_{k}(\widetilde{\theta})\Bigl{(}\nabla_{\widetilde{\theta}}\log\rho_{\widetilde{a}}^{\rm GM}(\widetilde{\theta})+\nabla_{\widetilde{\theta}}\widetilde{\Phi}_{R}(\widetilde{\theta})\Bigr{)}\mathrm{d}\widetilde{\theta},$
    $\dot{\widetilde{C}}_{k}=T\dot{C}_{k}T^{T}$ $=-TC_{k}\Bigr{(}\int\mathcal{N}_{k}(\theta)\bigl{(}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(\theta)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}(\theta)\bigr{)}\mathrm{d}\theta\Bigr{)}C_{k}T^{T}$
    $=-TC_{k}T^{T}\Bigr{(}\int\widetilde{\mathcal{N}}_{k}(\widetilde{\theta})T^{-T}\bigl{(}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}(\theta)+\nabla_{\theta}\nabla_{\theta}\Phi_{R}(\theta)\bigr{)}T^{-1}\mathrm{d}\widetilde{\theta}\Bigr{)}TC_{k}T^{T}$
    $=-\widetilde{C}_{k}\Bigl{(}\int\widetilde{\mathcal{N}}_{k}(\widetilde{\theta})\bigl{(}\nabla_{\widetilde{\theta}}\nabla_{\widetilde{\theta}}\log\rho_{\widetilde{a}}^{\rm GM}(\widetilde{\theta})+\nabla_{\widetilde{\theta}}\nabla_{\widetilde{\theta}}\widetilde{\Phi}_{R}(\widetilde{\theta})\bigr{)}\mathrm{d}\widetilde{\theta}\Bigr{)}\widetilde{C}_{k},$
    $\dot{\widetilde{w}}_{k}=\dot{w}_{k}$ $=-w_{k}\int\Bigl{(}\mathcal{N}_{k}(\theta)-\rho_{a}^{\rm GM}(\theta)\Bigr{)}\bigl{(}\log\rho_{a}^{\rm GM}(\theta)+\Phi_{R}(\theta)\bigr{)}\mathrm{d}\theta$
    $=-\widetilde{w}_{k}\int\Bigl{(}\widetilde{\mathcal{N}}_{k}(\widetilde{\theta})-\rho_{\widetilde{a}}^{\rm GM}(\widetilde{\theta})\Bigr{)}\bigl{(}\log|\det T|+\log\rho_{\widetilde{a}}^{\rm GM}(\widetilde{\theta})+\widetilde{\Phi}_{R}(\widetilde{\theta})\bigr{)}\mathrm{d}\widetilde{\theta}$
    $=-\widetilde{w}_{k}\int\Bigl{(}\widetilde{\mathcal{N}}_{k}(\widetilde{\theta})-\rho_{\widetilde{a}}^{\rm GM}(\widetilde{\theta})\Bigr{)}\bigl{(}\log\rho_{\widetilde{a}}^{\rm GM}(\widetilde{\theta})+\widetilde{\Phi}_{R}(\widetilde{\theta})\bigr{)}\mathrm{d}\widetilde{\theta},$
    where in the last equality, we used the fact that $\widetilde{\mathcal{N}}_{k}$ and $\rho_{\widetilde{a}}^{\rm GM}$ are both densities and thus $\int\bigl{(}\widetilde{\mathcal{N}}_{k}(\widetilde{\theta})-\rho_{\widetilde{a}}^{\rm GM}(\widetilde{\theta})\bigr{)}\log|\det T|\mathrm{d}\widetilde{\theta}=0$. This completes the proof.
  2. 2.
    To prove that the DF-GMVI algorithm Eq. 30 is $\mathcal{T}$-invariant, it suffices to show that the quadrature rules defined in Definition 3.1 and Definition 3.2 satisfy the following equations for any $T\in\mathcal{T}$:
    (35) $\mathrm{\mathsf{QR}}_{\widetilde{\mathcal{N}}_{k}}\bigl{\{}\nabla_{\widetilde{\theta}}\nabla_{\widetilde{\theta}}\log\rho_{\widetilde{a}}^{\rm GM}+\nabla_{\widetilde{\theta}}\nabla_{\widetilde{\theta}}\widetilde{\Phi}_{R}\bigr{\}}$ $=T^{-T}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{\}}T^{-1},$
    $\mathrm{\mathsf{QR}}_{\widetilde{\mathcal{N}}_{k}}\bigl{\{}\nabla_{\widetilde{\theta}}\log\rho_{\widetilde{a}}^{\rm GM}+\nabla_{\widetilde{\theta}}\widetilde{\Phi}_{R}\bigr{\}}$ $=T^{-T}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\nabla_{\theta}\log\rho_{a}^{\rm GM}+\nabla_{\theta}\Phi_{R}\bigr{\}},$
    $\mathrm{\mathsf{QR}}_{\widetilde{\mathcal{N}}_{k}}\bigl{\{}\log\rho_{\widetilde{a}}^{\rm GM}+\widetilde{\Phi}_{R}\bigr{\}}$ $=\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\bigl{\{}\log\rho_{a}^{\rm GM}+\Phi_{R}\bigr{\}}-\log|\det T|.$
    Here in the last equality, we can tolerate a constant difference since we will normalize ${w_{k}(t+\Delta t)}_{k=1}^{K}$ after the update Eq. 30. Given any Gaussian density $\mathcal{N}=\mathcal{N}(\theta;m,C)$, an affine mapping transforms it to another Gaussian $\widetilde{\mathcal{N}}=\mathcal{N}(\widetilde{\theta};\widetilde{m},\widetilde{C})$. The square root matrix computed by Cholesky decomposition satisfies $\sqrt{\widetilde{C}}=\sqrt{TCT^{T}}=T\sqrt{C}$, when $T$ is an invertible lower triangular matrix. Hence, the quadrature points generated for the two Gaussian densities according to Definition 3.1 satisfy
    $\widetilde{\theta}_{0}$ $=\widetilde{m}=Tm+d=T\theta_{0}+d,$
    $\widetilde{\theta}_{i}$ $=\widetilde{m}+\alpha[\sqrt{\widetilde{C}}]_{i}=Tm+d+\alpha T[\sqrt{C}]_{i}=T\theta_{i}+d,$
    $\widetilde{\theta}_{N_{\theta}+i}$ $=\widetilde{m}-\alpha[\sqrt{\widetilde{C}}]_{i}=Tm+d-\alpha T[\sqrt{C}]_{i}=T\theta_{N_{\theta}+i}+d,$
    for $1\leq i\leq N_{\theta}$. These equations imply that ~$\widetilde{\mathcal{F}}(\widetilde{\theta_{i}})=\mathcal{F}(\theta_{i})$ for $0\leq i\leq 2N_{\theta}$, and thus $\widetilde{c}=c$, $\widetilde{b_{i}}=b_{i}$, $\widetilde{a_{i}}=a_{i}$ for $1\leq i\leq N_{\theta}$, and then $\widetilde{B}=B$, $\widetilde{A}=A$. Consequently,
    (36) $\mathrm{\mathsf{QR}}_{\widetilde{\mathcal{N}}}\{\widetilde{\Phi}_{R}\}$ $=\frac{1}{2}\widetilde{c}^{T}\widetilde{c}=\frac{1}{2}c^{T}c=\mathrm{\mathsf{QR}}_{\mathcal{N}}\{\Phi_{R}\},$
    $\mathrm{\mathsf{QR}}_{\widetilde{\mathcal{N}}}\{\nabla_{\widetilde{\theta}}\widetilde{\Phi}_{R}\}$ $=\sqrt{\widetilde{C}}^{-T}\widetilde{B}^{T}\widetilde{c}=T^{-T}\sqrt{C}^{-T}B^{T}C=T^{-T}\mathrm{\mathsf{QR}}_{\mathcal{N}}\{\nabla_{\theta}\Phi_{R}\},$
    $\mathrm{\mathsf{QR}}_{\widetilde{\mathcal{N}}}\{\nabla_{\widetilde{\theta}}\nabla_{\widetilde{\theta}}\widetilde{\Phi}_{R}\}$ $=\sqrt{\widetilde{C}}^{-T}(6\mathrm{Diag}(\widetilde{A}^{T}\widetilde{A})+\widetilde{B}^{T}\widetilde{B})\sqrt{\widetilde{C}}^{-1}$
    $=T^{-T}\sqrt{C}^{-T}(6\mathrm{Diag}(A^{T}A)+B^{T}B)\sqrt{C}^{-1}T^{-1}$
    $=T^{-T}\mathrm{\mathsf{QR}}_{\mathcal{N}}\{\nabla_{\theta}\nabla_{\theta}\Phi_{R}\}T^{-1}.$
    By the Definition 3.2, we have that
    (37) $\mathrm{\mathsf{QR}}_{\widetilde{\mathcal{N}}_{k}}\{\log\rho_{\widetilde{a}}^{\rm GM}\}=\log\rho_{\widetilde{a}}^{\rm GM}(\widetilde{m}_{k})=\log\left(|\det T|^{-1}\rho_{a}^{\rm GM}(m_{k})\right)=\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\{\log\rho_{a}^{\rm GM}\}-\log|\det T|,$
    $\mathrm{\mathsf{QR}}_{\widetilde{\mathcal{N}}_{k}}\{\nabla_{\widetilde{\theta}}\log\rho_{\widetilde{a}}^{\rm GM}\}=\nabla_{\widetilde{\theta}}\log\rho_{\widetilde{a}}^{\rm GM}(\widetilde{m}_{k})=T^{-T}\nabla_{\theta}\log\rho_{a}^{\rm GM}(m_{k})=T^{-T}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\{\nabla_{\theta}\log\rho_{a}^{\rm GM}\}.$
    Using the same notations as in Definition 3.2, we have $\widetilde{v_{i}}(\widetilde{\theta})=\widetilde{C}_{i}^{-1}(\widetilde{\theta}-\widetilde{m_{i}})=(TC_{i}T^{T})^{-1}(T\theta-Tm_{i})=T^{-T}v_{i}(\theta)$, and thus it is easy to check that
    (38) $\mathrm{\mathsf{QR}}_{\widetilde{\mathcal{N}_{k}}}\{\nabla_{\widetilde{\theta}}\nabla_{\widetilde{\theta}}\log\rho_{\widetilde{a}}^{\rm GM}\}=T^{-T}\mathrm{\mathsf{QR}}_{\mathcal{N}_{k}}\{\nabla_{\theta}\nabla_{\theta}\log\rho_{a}^{\rm GM}\}T^{-1}.$
    Combining Eq. 36, Eq. 37 and Eq. 38, we finish the proof of Eq. 35.